Two particles A and B are projected simultaneously from the
two towers of height 10 m and 20 m respectively. Particle A is
projected with an initial speed of 10√2 m/s at an angle of 45
with horizontal, while particle B is projected horizontally with
speed 10 m/s. If they collide in air, what is the distance d
between the towers?

Grade:

2 Answers

Khimraj
4 years ago
No they will not collide as vertical component is not same for both projectiles.
Khimraj
4 years ago
Vo1 = 10√2 
Vo1y = Vo1x = 10√2*√2/2 = 10.0 m8sec 
Vo2x = 10m/sec 

if they meet, then : 
h1+Vo1y*t-4.9t^2 = h2-4.9t^2 
10+10t = 20+4.9t^2-4.9t^2 
10t = 10 
t = 1 
distance = (Vo1x+Vo2x)*t = (10+10)*1 = 20.0 m